OpenStudy (anonymous):

If you are finding the line perpendicular to x + 3y = 5, what is the slope of the line?

OpenStudy (anonymous):

3y=-x+5 y= -1/3x+5 slope is -1/3

OpenStudy (aroub):

3y=-x+5 y=-1/3x+5/3 slope = 3

OpenStudy (anonymous):

ohh for got perpendiculer is negative reciprocal so it 3

OpenStudy (anonymous):

3